OpenStudy (anonymous):

2aQ= 6xQ/5 +b ....how do you solve for Q????

OpenStudy (anonymous):

2aQ = 6xQ/5 + b Divide by Q: 2a = 6x/5 + b/Q Move the term without Q so b/Q is by itself: 2a - 6x/5 = b/Q Remove the b so Q is by itself: 2a/b - 6x/5b = 1/Q Invert to get Q: 5b/6x - b/2a = Q

OpenStudy (anonymous):

what happened to the other Q though...there were 2 in the original equation..which is where im confues

OpenStudy (anonymous):

Dividing both sides by Q gives 2aQ/Q = 6xQ/5Q + b/Q We can then cancel the Qs from the 2a term and the 6x term keeping the one gained by the + b term. This leaves 2a = 6x/5 + b/Q One Q is what we want.